OpenStudy (anonymous):

Hi, does anyone know which mathematics courses I should complete before starting this course? Is single variable calculus sufficient?

OpenStudy (waynex):

Single variable is all you need. It can be completed concurrently with this course. The professor does mention a couple differential equations during lecture, but it's not something that is covered in the tests.

OpenStudy (zenharmony):

There is a link in the Prerequisites section of the Course Detail page with which you can access the Single Variable Calculus course, as well.

OpenStudy (jh3power):

\[\frac{ dx }{ dt }= \lim_{t \rightarrow 0}\frac{ \Delta x }{ \Delta t }\]\[ \frac{ dx^n }{ dt } = nx^{n-1}\] and the definition of \[\int\limits_{0}^{?}\] should matter. You don't need much, I think.Only that...